# Abelian congruences and similarity in varieties with a weak difference term

### v1

@@ 7. Similarity
In a 1983 paper [5], Freese introduced a relation of similarity between subdirectly irreducible algebras in a congruence modular variety. If the monoliths of the algebras are nonabelian, then the algebras are similar if and only if they are isomorphic. If the monoliths are abelian, then the relation is more interesting. Freese’s original definition in the latter case, as well as the equivalent formulation found in [6], depend on a certain “uniformity” of the blocks of an abelian monolith $\mu$ contained in a common $(0:\mu)$-block, represented in part by Corollary 5.3. As we have claimed (and will prove in Section 8), this uniformity is lost in varieties with a weak difference term.
@@
    If $\mu$ is abelian, then $D({\mathbf{A}})=D({\mathbf{A}},\mu)$ as defined in Definition 4.7.
@@
Before stating the main theorem of this section, we introduce a variant of a crucial construct due to Zhuk [37, Section 6.6].
@@
Suppose ${\mathbf{A}},{\mathbf{B}}$ are subdirectly irreducible algebras in a common signature with monoliths $\mu,\kappa$ respectively. A proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$ is a subuniverse $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ satisfying
@@
If $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$, then we define its trace to be the set $\operatorname{tr}(T)=\{(a,b)\in A\times B\,:\,\text{$(a,a,b,b)\in T$}\}$, and its kernel to be the set
@@
Observe that for any proper bridge $T$ we have $\operatorname{tr}(T)=\operatorname{pr}_{1,3}(T)=\operatorname{pr}_{2,4}(T)$ by property (B3), and hence $\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ by property (B1).
In the special case where one of the two monoliths in the previous definition is nonabelian, the existence of a proper bridge is possible only in the most trivial way.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with monoliths $\mu,\kappa$ respectively, and $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Assume that one of $\mu$ or $\kappa$ is nonabelian. Then so is the other, $\operatorname{tr}(T)$ is the graph of an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$, and $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$.
@@
One can easily show, using (B1) and the definitions of $S$ and $Y$, that $X(\kappa,S)\subseteq Y$ (see Definition 2.1). Since $Y$ is a subuniverse of ${\mathbf{B}}^{2\times 2}$, we get $M(\kappa,S)\subseteq Y$. Then from (B2) and Definition 2.2(2) we can deduce $C(S,\kappa;0_{B})$. Since $S$ is a tolerance, we can improve this to $C(\operatorname{Cg}(S),\kappa;0_{B})$ by [17, Theorem 2.19(2)]. Now assume that $S\neq 0_{B}$; then $\operatorname{Cg}(S)\geq\kappa$ since $\kappa$ is the monolith of ${\mathbf{B}}$; but then $\kappa$ would be abelian, contrary to our assumption. This proves $S=0_{B}$ and hence $\operatorname{tr}(T)$ is the graph of a surjective homomorphism $h:{\mathbf{A}}\rightarrow{\mathbf{B}}$.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly directly algebras with monoliths $\mu,\kappa$ respectively, and $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Let ${\mathbf{C}}=\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ and $\tau=\ker(T)$. Then $\tau\in\operatorname{Con}({\mathbf{C}})$. Hence
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly directly, and $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Let ${\mathbf{C}}=\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ and $\tau=\ker(T)$. Also let $\delta_{1},\delta_{2}\in\operatorname{Con}({\mathbf{C}})$ be the kernels of the two projection homomorphisms $\operatorname{pr}_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $\operatorname{pr}_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and let $\delta^{+}_{i}$ be the unique upper cover of $\delta_{i}$ in $\operatorname{Con}({\mathbf{C}})$ for $i=1,2$. Then $(\delta_{1},\delta^{+}_{1})\searrow(0,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$.
@@
Finally, suppose $((a_{1},b_{1}),(a_{2},b_{2}))\in\tau\cap\delta_{1}$. Then $a_{1}=a_{2}$, so $b_{1}=b_{2}$ by (B2), so $(a_{1},b_{1})=(a_{2},b_{2})$. This shows $\tau\wedge\delta_{1}=0$, and a similar proof shows $\tau\wedge\delta_{2}=0$. This proves $(\delta_{1},\delta^{+}_{1})\searrow(0,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$ as required. ∎
In the next lemma, we show that proper bridges between subdirectly irreducible algebras with abelian monoliths satisfy additional properties.
Lemma 7.7.
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly directly algebras with abelian monoliths $\mu,\kappa$ respectively, and $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Let $\alpha=(0:\mu)$ and $\beta=(0:\kappa)$; also let $\Delta_{{\mathbf{A}}}=\Delta_{\mu,\alpha}$ and $\Delta_{{\mathbf{B}}}=\Delta_{\kappa,\beta}$.
  1. (1)
    For all $(a,r),(b,s)\in\operatorname{tr}(T)$, $(a,b)\in\alpha\iff(r,s)\in\beta$.
  2. (2)
    For every $(\mu,\alpha)$-matrix $M=\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}$ there exists a $(\kappa,\beta)$-matrix $\begin{pmatrix}r&r^{\prime}\\ s&s^{\prime}\end{pmatrix}$ with $(a,b,r,s),(a^{\prime},b^{\prime},r^{\prime},s^{\prime})\in T$.
  3. (3)
    For all $(a,b,r,s),(a^{\prime},b^{\prime},r^{\prime},s^{\prime})\in T$, $(a,b)\stackrel{{\scriptstyle\Delta_{{\mathbf{A}}}}}{{\equiv}}(a^{\prime},b^{\prime})\iff(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r^{\prime},s^{\prime})$.
Proof.
(1) Define ${\mathbf{C}}=\operatorname{tr}(T)\leq{\mathbf{A}}\times{\mathbf{B}}$ and $\delta_{1},\delta^{+}_{1},\delta_{2},\delta^{+}_{2},\tau\in\operatorname{Con}({\mathbf{C}})$ as in Lemma 7.5. Observe that $\delta^{+}_{i}/\delta_{i}$ is abelian for $i=1,2$, because $\mu$ and $\kappa$ are abelian. Since $(\delta_{1},\delta^{+}_{1})\searrow(0,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$ by Lemma 7.5, we get $(\delta_{1}:\delta^{+}_{1})=(0:\tau)=(\delta_{2}:\delta^{+}_{2})$ by Proposition 3.13(1). On the other hand, letting $\operatorname{pr}_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ be the first projection homomorphism, we have $(\delta_{1}:\delta^{+}_{1})=(\operatorname{pr}_{1}^{-1}(0):\operatorname{pr}_{1}^{-1}(\mu))=\operatorname{pr}_{1}^{-1}((0:\mu))=\operatorname{pr}_{1}^{-1}(\alpha)$ by Lemma 2.3(2), and similarly $(\delta_{2}:\delta^{+}_{2})=\operatorname{pr}_{2}^{-1}(\beta)$. Together these facts prove $\operatorname{pr}_{1}^{-1}(\alpha)=\operatorname{pr}_{2}^{-1}(\beta)$, which is equivalent to (1).
(2) Recall that the set $M(\mu,\alpha)$ of all $(\mu,\alpha)$-matrices is the subuniverse of ${\mathbf{A}}^{2\times 2}$ generated by the set
$X(\mu,\alpha):=\left\{\begin{pmatrix}a&a\\ b&b\end{pmatrix}:(a,b)\in\mu\right\}\cup\left\{\begin{pmatrix}a&a^{\prime}\\ a&a^{\prime}\end{pmatrix}:(a,a^{\prime})\in\alpha\right\}.$
Let $S$ be the set of pairs $(M,N)=\left(\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix},\begin{pmatrix}r&r^{\prime}\\ s&s^{\prime}\end{pmatrix}\right)$ where $M$ is a $(\mu,\alpha)$-matrix, $N$ is a $(\kappa,\beta)$-matrix, and $(a,b,r,s),(a^{\prime},b^{\prime},r^{\prime},s^{\prime})\in T$. Then $S\leq{\mathbf{A}}^{2\times 2}\times{\mathbf{B}}^{2\times 2}$ and $\operatorname{pr}_{1}(S)$ is a subuniverse of $M(\mu,\alpha)$. As we are desiring to prove $\operatorname{pr}_{1}(S)=M(\mu,\alpha)$, it is therefore enough to prove $X(\mu,\alpha)\subseteq\operatorname{pr}_{1}(S)$. That is, to prove (2), it is enough to prove it for matrices $M\in X(\mu,\alpha)$.
Consider first the case where $M=\begin{pmatrix}a&a\\ b&b\end{pmatrix}$ with $(a,b)\in\mu$. By (B1), there exists $(r,s)\in\kappa$ with $(a,b,r,s)\in T$. Then $N:=\begin{pmatrix}r&r\\ s&s\end{pmatrix}\in M(\kappa,\beta)$ and $(M,N)\in S$ witnessed by $(a,b,r,s)$ and $(a^{\prime},b^{\prime},r^{\prime},s^{\prime}):=(a,b,r,s)$.
Consider next the case where $M=\begin{pmatrix}a&a^{\prime}\\ a&a^{\prime}\end{pmatrix}$ with $(a,a^{\prime})\in\alpha$. Pick $r,r^{\prime}\in B$ with $(a,r),(a^{\prime},r^{\prime})\in\operatorname{tr}(T)$. Then $(r,r^{\prime})\in\beta$ by (1), so $N:=\begin{pmatrix}r&r^{\prime}\\ r&r^{\prime}\end{pmatrix}\in M(\kappa,\beta)$, and we have $(M,N)\in S$ witnessed by $(a,a,r,r)$ and $(a^{\prime},a^{\prime},r^{\prime},r^{\prime})$. This completes the proof of (2).
We start the proof of (3) by proving the following special case.
Claim 7.7.1.
For all $(a,b)\in\mu$ and $(r,s),(u,v)\in\kappa$,
$(a,b,r,s),(a,b,u,v)\in T\implies(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(u,v).$
Proof of Claim 7.7.1.
Assume $(a,b,r,s),(a,b,u,v)\in T$. Then $(a,r),(a,u)\in\operatorname{tr}(T)$, so $(r,u)\in\beta$ by (1). Since $(r,s),(u,v)\in\kappa\leq\beta$, it follows that $r,s,u,v$ belong to a common $\beta$-class.
Let $d(x,y,z)$ be a weak difference term for $\mathcal{V}$. We claim that
($\ast$) $d(s,u,u)=d(s,v,v)\quad\mbox{and}\quad d(r,u,u)=d(s,v,u).$
To prove the first equality, observe that $d(\underline{u},u,u)=u=d(\underline{u},v,v)$ since $u,v$ belong to a common class of an abelian congruence ($\kappa$). Since $(u,s)\in\beta$ and $C(\beta,\kappa;0)$, we can replace the underlined occurrences of $u$ with $s$ to get $d(\underline{s},u,u)=d(\underline{s},v,v)$.
To prove the second equality, observe that the following three 4-tuples belong to $T$ (the third because of (B3)):
$(a,b,r,s),\quad(a,b,u,v),\quad(a,a,u,u).$
Applying the weak difference term coordinate-wise to these tuples gives
$(a,d(b,b,a),d(r,u,u),d(s,v,u))\in T.$
As $a,b$ belong to a common class of an abelian congruence ($\mu$), we get $d(b,b,a)=a$. Thus by bridge property (B2) we get $d(r,u,u)=d(s,v,u)$.
Now we build some $(\kappa,\beta)$-matrices. Let
$M_{1}=\begin{pmatrix}r&r\\ s&s\end{pmatrix},~{}~{}M_{2}=\begin{pmatrix}r&u\\ r&u\end{pmatrix},~{}~{}M_{3}=\begin{pmatrix}s&v\\ s&v\end{pmatrix},~{}~{}M_{4}=\begin{pmatrix}v&v\\ v&v\end{pmatrix},~{}~{}M_{5}=\begin{pmatrix}u&u\\ v&v\end{pmatrix}.$
Clearly $M_{1},\ldots,M_{5}\in M(\kappa,\beta)$. Applying the weak difference term gives
$N:=d(M_{1},M_{2},M_{2})$ $=\begin{pmatrix}r&d(r,u,u)\\ s&d(s,u,u)\end{pmatrix}\in M(\kappa,\beta)$
$N^{\prime}:=d(M_{3},M_{4},M_{5})$ $=\begin{pmatrix}d(s,v,u)&u\\ d(s,v,v)&v\end{pmatrix}\in M(\kappa,\beta).$
The second column of $N$ and the first column of $N^{\prime}$ are identical by ($\ast$7). Since $\Delta_{{\mathbf{B}}}$ is the “horizontal transitive closure” of $M(\kappa,\beta)$, it follows that $(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(u,v)$ as required. ∎
Now we prove (3). Fix $(a,b,r,s),(a^{\prime},b^{\prime},r^{\prime},s^{\prime})\in T$. Assume $(a,b)\stackrel{{\scriptstyle\Delta_{{\mathbf{A}}}}}{{\equiv}}(a^{\prime},b^{\prime})$. Then there exist matrices
$\begin{pmatrix}a_{0}&a_{1}\\ b_{0}&b_{1}\end{pmatrix},\begin{pmatrix}a_{1}&a_{2}\\ b_{1}&b_{2}\end{pmatrix},\ldots,\begin{pmatrix}a_{n-1}&a_{n}\\ b_{n-1}&b_{n}\end{pmatrix}\in M(\mu,\alpha)$
with $(a_{0},b_{0})=(a,b)$ and $(a_{n},b_{n})=(a^{\prime},b^{\prime})$. For each $i<n$, apply item (2) to the matrix $\begin{pmatrix}a_{i}&a_{i+1}\\ b_{i}&b_{i+1}\end{pmatrix}$ to get a matrix
$\begin{pmatrix}r_{i}&r_{i}^{\prime}\\ s_{i}&s_{i}^{\prime}\end{pmatrix}\in M(\kappa,\beta)$
with
$(a_{i},b_{i},r_{i},s_{i})\in T\quad\mbox{and}\quad(a_{i+1},b_{i+1},r_{i}^{\prime},s_{i}^{\prime})\in T.$
Then for each $i<n$ we have $(r_{i},s_{i})\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r_{i}^{\prime},s_{i}^{\prime})$ since these pairs are the two columns of a $(\kappa,\beta)$-matrix and $\Delta_{{\mathbf{B}}}$ is the transitive closure of $M_{\kappa,\beta}$. By Claim 7.7.1 we also get
$(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r_{0},s_{0}),\quad\mbox{$(r_{i}^{\prime},s_{i}^{\prime})\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r_{i+1},s_{i+1})$ for all $i<n$},\quad\mbox{and}\quad(r_{n}^{\prime},s_{n}^{\prime})\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r^{\prime},s^{\prime}).$
Then $(r,s)\stackrel{{\scriptstyle\Delta_{{\mathbf{B}}}}}{{\equiv}}(r^{\prime},s^{\prime})$ by transitivity. The converse implication is proved similarly. ∎
@@
    There exist an algebra ${\mathbf{C}}\in\mathcal{V}$, surjective homomorphisms $f_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and congruences $\psi,\tau\in\operatorname{Con}({\mathbf{C}})$ with $\psi<\tau$, such that, letting $\delta_{i}=\ker(f_{i})$ and letting $\delta^{+}_{i}$ denote the unique upper cover of $\delta_{i}$ in $\operatorname{Con}({\mathbf{C}})$ for $i=1,2$, we have $(\delta_{1},\delta^{+}_{1})\searrow(\psi,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$.
@@
    There exists a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$.
@@
Throughout this proof, let $\mu,\kappa$ denote the monoliths of ${\mathbf{A}},{\mathbf{B}}$ respectively. The implication (3) $\Rightarrow$ (2) is clear, and (4) $\Rightarrow$ (3) holds by Lemma 7.6. We will first prove (2) $\Rightarrow$ (1) $\Rightarrow$ (4) in the case when one of $\mu$ or $\kappa$ is nonabelian. Then we will prove (2) $\Rightarrow$ (4) $\Leftrightarrow$ (1) in the case when both $\mu$ and $\kappa$ are abelian.
@@
(1) $\Rightarrow$ (4). Since ${\mathbf{A}}\sim{\mathbf{B}}$, it follows that both $\mu$ and $\kappa$ are nonabelian and there exists an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$. Then the set $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$.
(2) $\Rightarrow$ (1). Assume there exist ${\mathbf{C}}\in\mathcal{V}$, surjective homomorphisms $f_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and congruences $\psi,\tau\in\operatorname{Con}({\mathbf{C}})$ with $\psi<\tau$, such that, letting $\delta_{i}=\ker(f_{i})$ and letting $\delta^{+}_{i}$ denote the unique upper cover of $\delta_{i}$ for $i=1,2$, we have $(\delta_{1},\delta^{+}_{1})\searrow(\psi,\tau)\nearrow(\delta_{2},\delta^{+}_{2})$. Because we are in Case 1, at least one of $\delta^{+}_{1}/\delta_{1}$ or $\delta^{+}_{2}/\delta_{2}$ is nonabelian. By Proposition 3.12(1), both are nonabelian, as is $\tau/\psi$. It follows that $\mu$ and $\kappa$ are both nonabelian.
For each $i=1,2$ we have $\delta_{i}\wedge\tau=\psi$ and hence $C(\delta_{i},\tau;\psi)$; see e.g. [17, Theorem 2.19(8)]; thus $C(\delta_{1}\vee\delta_{2},\tau;\psi)$ by semidistributivity of the centralizer relation in the first variable; see e.g. [17, Theorem 2.19(5)]. Suppose $\delta_{1}\neq\delta_{2}$; then without loss of generality we can assume $\delta_{1}\vee\delta_{2}>\delta_{2}$. Hence $\delta_{1}\vee\delta_{2}\geq\delta^{+}_{2}$ (as $\delta_{2}$ is completely meet-irreducible) and thus $\delta_{1}\vee\delta_{2}\geq\tau$. Then $C(\delta_{1}\vee\delta_{2},\tau;\psi)$ implies $C(\tau,\tau;\psi)$ by monotonicity, which contradicts the fact that $\tau/\psi$ is nonabelian. This proves $\delta_{1}=\delta_{2}$ and hence ${\mathbf{A}}\cong{\mathbf{C}}/\delta_{1}={\mathbf{C}}/\delta_{2}\cong{\mathbf{B}}$. Hence ${\mathbf{A}}\sim{\mathbf{B}}$.
@@
(4) $\Rightarrow$ (1). Assume that $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. By Lemma 7.7(3), the set
$\{((a,b)/\Delta_{{\mathbf{A}}},(r,s)/\Delta_{{\mathbf{B}}})\,:\,\text{$(a,b,r,s)\in T$}\}$
is the graph of a bijection $\lambda$. By the nature of its definition it is also a subalgebra of $D({\mathbf{A}})\times D({\mathbf{B}})$, and by (B1) the domain and range of $\lambda$ are $D({\mathbf{A}})$ and $D({\mathbf{B}})$ respectively. Thus $\lambda$ is an isomorphism from $D({\mathbf{A}})$ to $D({\mathbf{B}})$.
(1) $\Rightarrow$ (4). Assume $\lambda_{0}:D({\mathbf{A}})\cong D({\mathbf{B}})$. Recall that $D({\mathbf{A}})$ and $D({\mathbf{B}})$ are subdirectly irreducible with monoliths $\varphi_{{\mathbf{A}}},\varphi_{{\mathbf{B}}}$ respectively, by Corollary 4.8. Thus we must have $\lambda_{0}(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$. As $D^{o}_{{\mathbf{A}}}$ is a subuniverse of $D({\mathbf{A}})$ and a transversal for $\varphi_{{\mathbf{A}}}$, it follows that $\lambda_{0}(D^{o}_{{\mathbf{A}}})$, like $D^{o}_{{\mathbf{B}}}$, is a subuniverse of $D({\mathbf{B}})$ and a transversal for $\varphi_{{\mathbf{B}}}$. Thus by Lemma 3.9, there exists an isomorphism $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$ satisfying $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$. Now define
[…]
@@
Again suppose $(a_{1},a_{2},b_{1},b_{2})\in T$ and let $r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})$ and $s=(b_{1},b_{1})/\Delta_{{\mathbf{B}}}$. Observe that $(a_{1},a_{1})/\Delta_{{\mathbf{A}}}\stackrel{{\scriptstyle\varphi_{{\mathbf{A}}}}}{{\equiv}}(a_{1},a_{2})/\Delta_{{\mathbf{A}}}$ and $(b_{1},b_{1})/\Delta_{{\mathbf{B}}}\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}(b_{1},b_{2})/\Delta_{{\mathbf{B}}}$. By an earlier observation we have $\lambda(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$. Thus
$r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})$ $\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}\lambda((a_{1},a_{2})/\Delta_{{\mathbf{A}}})$

### v2

@@ 7. Similarity
In his 1983 paper [5], Freese introduced a relation of similarity between subdirectly irreducible algebras in a congruence modular variety. If the monoliths of the algebras are nonabelian, then the algebras are similar if and only if they are isomorphic. If the monoliths are abelian, then the relation is more interesting. Freese’s original definition in the latter case, as well as the equivalent formulation found in [6], relied heavily on having a difference term, and does not easily adapt to varieties with a weak difference term.
@@
    If $\mu$ is abelian, then $D({\mathbf{A}})$ is $D({\mathbf{A}},\mu)$, the difference algebra for $\mu$.
@@
Before stating the main theorem of this section, we introduce a variant of a crucial construct due to Zhuk [35, Section 6.6].
@@
Suppose ${\mathbf{A}},{\mathbf{B}}$ are subdirectly irreducible algebras in a common signature with monoliths $\mu,\kappa$ respectively. A similarity bridge from $\mathbf{A}$} to $\mathbf{B}$} is a subuniverse $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$ satisfying
@@
If $T$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}, then we define its trace to be the set $\operatorname{tr}(T)=\{(a,b)\in A\times B\,:\,\text{$(a,a,b,b)\in T$}\}$, and its kernel to be the set
@@
Observe that for any similarity bridge $T$ we have $\operatorname{tr}(T)=\operatorname{pr}_{1,3}(T)=\operatorname{pr}_{2,4}(T)$ by property (B3), and hence $\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ by property (B1).
In the special case where one of the two monoliths in the previous definition is nonabelian, the existence of a similarity bridge is possible only in the most trivial way.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with monoliths $\mu,\kappa$ respectively, and $T$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}. Assume that one of $\mu$ or $\kappa$ is nonabelian. Then so is the other, $\operatorname{tr}(T)$ is the graph of an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$, and $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$.
@@
One can easily show, using (B1) and the definitions of $S$ and $Y$, that $X(\kappa,S)\subseteq Y$ (see Definition 2.1). Since $Y$ is a subuniverse of ${\mathbf{B}}^{2\times 2}$, we get $M(\kappa,S)\subseteq Y$. Then from (B2) and Definition 2.2(2) we can deduce $C(S,\kappa;0_{B})$. Since $S$ is a tolerance, we can improve this to $C(\operatorname{Cg}(S),\kappa;0_{B})$ by [16, Theorem 2.19(2)]. Now assume that $S\neq 0_{B}$; then $\operatorname{Cg}(S)\geq\kappa$ since $\kappa$ is the monolith of $\mathbf{B}$}; but then $\kappa$ would be abelian, contrary to our assumption. This proves $S=0_{B}$ and hence $\operatorname{tr}(T)$ is the graph of a surjective homomorphism $h:{\mathbf{A}}\rightarrow{\mathbf{B}}$.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly directly algebras with monoliths $\mu,\kappa$ respectively, and $T$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}. Let ${\mathbf{C}}=\operatorname{tr}(T)\leq_{sd}{\mathbf{A}}\times{\mathbf{B}}$ and $\tau=\ker(T)$. Then $\tau\in\operatorname{Con}({\mathbf{C}})$. Hence
@@
[…]
@@
[…]
@@
[…]
@@
    There exists a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}.
@@
Throughout this proof, let $\mu,\kappa$ denote the monoliths of ${\mathbf{A}},{\mathbf{B}}$ respectively. The implication (3) $\Rightarrow$ (2) is clear, and (4) $\Rightarrow$ (3) holds by Lemma 7.6. We will prove (2) $\Rightarrow$ (1) $\Rightarrow$ (4) first in the case when one of $\mu$ or $\kappa$ is nonabelian, and then in the case when both $\mu$ and $\kappa$ are abelian.
@@
(1) $\Rightarrow$ (4). Since ${\mathbf{A}}\sim{\mathbf{B}}$, it follows that both $\mu$ and $\kappa$ are nonabelian and there exists an isomorphism $h:{\mathbf{A}}\cong{\mathbf{B}}$. Then the set $T=\{(a,b,h(a),h(b))\,:\,\text{$(a,b)\in\mu$}\}$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}.
[…]
For each $i=1,2$ we have $\delta_{i}\wedge\tau=\psi$ and hence $C(\delta_{i},\tau;\psi)$; see e.g. [16, Theorem 2.19(8)]; thus $C(\delta_{1}\vee\delta_{2},\tau;\psi)$ by semidistributivity of the centralizer relation in the first variable; see e.g. [16, Theorem 2.19(5)]. Suppose $\delta_{1}\neq\delta_{2}$; then without loss of generality we can assume $\delta_{1}\vee\delta_{2}>\delta_{2}$. Hence $\delta_{1}\vee\delta_{2}\geq\delta^{+}_{2}$ (as $\delta_{2}$ is completely meet-irreducible) and thus $\delta_{1}\vee\delta_{2}\geq\tau$. Then $C(\delta_{1}\vee\delta_{2},\tau;\psi)$ implies $C(\tau,\tau;\psi)$ by monotonicity, which contradicts the fact that $\tau/\psi$ is nonabelian. This proves $\delta_{1}=\delta_{2}$ and hence ${\mathbf{A}}\cong{\mathbf{C}}/\delta_{1}={\mathbf{C}}/\delta_{2}\cong{\mathbf{B}}$. Hence ${\mathbf{A}}\sim{\mathbf{B}}$.
@@
(2) $\Rightarrow$ (1) We have $D({\mathbf{A}})\cong D({\mathbf{C}}/\delta_{1},\delta_{1}^{+}/\delta_{1})$ and $D({\mathbf{B}})\cong D({\mathbf{C}}/\delta_{2},\delta_{2}^{+}/\delta_{2})$ since ${\mathbf{A}}\cong{\mathbf{C}}/\delta_{1}$ and ${\mathbf{B}}\cong{\mathbf{C}}/\delta_{2}$. We can replace $\tau$ with any $\tau^{\prime}\in\operatorname{Con}({\mathbf{C}})$ satisfying $\psi\prec\tau^{\prime}\leq\tau$ and the hypotheses of (2) still hold; thus we may assume $\psi\prec\tau$. We have that $\tau/\psi$ is abelian, since $(\psi,\tau)\nearrow(\delta_{1},\delta_{1}^{+})$ and $\delta_{1}^{+}/\delta_{1}$ is abelian. Then by Theorem 6.10 we get $D({\mathbf{C}}/\delta_{1},\delta_{1}^{+}/\delta_{1})\cong D({\mathbf{C}}/\psi,\tau/\psi)\cong D({\mathbf{C}}/\delta_{2},\delta_{2}^{+}/\delta_{2})$. Hence $D({\mathbf{A}})\cong D({\mathbf{B}})$, so ${\mathbf{A}}\sim{\mathbf{B}}$.
(1) $\Rightarrow$ (4). Assume $D({\mathbf{A}})\cong D({\mathbf{B}})$. As in the proof of Proposition 6.9, there exists an isomorphism $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$ satisfying $\lambda(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$ and $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$. Define
$T=\{(a_{1},a_{2},b_{1},b_{2})\in A\times A\times B\times B$ $:~\mbox{$a_{1}\stackrel{{\scriptstyle\mu}}{{\equiv}}a_{2}$ and $b_{1}\stackrel{{\scriptstyle\kappa}}{{\equiv}}b_{2}$ and}$
@@
Again suppose $(a_{1},a_{2},b_{1},b_{2})\in T$ and let $r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})$ and $s=(b_{1},b_{1})/\Delta_{{\mathbf{B}}}$. Observe that $(a_{1},a_{1})/\Delta_{{\mathbf{A}}}\stackrel{{\scriptstyle\varphi_{{\mathbf{A}}}}}{{\equiv}}(a_{1},a_{2})/\Delta_{{\mathbf{A}}}$ and $(b_{1},b_{1})/\Delta_{{\mathbf{B}}}\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}(b_{1},b_{2})/\Delta_{{\mathbf{B}}}$. Then
$r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})$ $\stackrel{{\scriptstyle\varphi_{{\mathbf{B}}}}}{{\equiv}}\lambda((a_{1},a_{2})/\Delta_{{\mathbf{A}}})$ as $\lambda(\varphi_{{\mathbf{A}}})=\varphi_{{\mathbf{B}}}$